A 50kg student walks and runs up a set of stairs that has a vertical height of 1.5meters. It takes 3 seconds to walk up and 1 second to run up. Compare the work and power done walking and running.


1
Expert's answer
2021-05-09T13:04:31-0400

"\frac{A_r}{A_w}=\frac{mgh}{mgh}=1,"

"\frac{P_r}{P_w}=\frac{A_r}{t_r}\cdot \frac{t_w}{A_w}=\frac{t_w}{t_r}=\frac 31=3."